OpenStudy (anonymous):

Jules kicks a soccer ball off the ground and into the air with an initial velocity of 25 feet per second. Assume the starting height of the ball is 0 feet. Approximately, how long does it take until the soccer ball hits the ground again?

OpenStudy (migitmack):

oh man, I hate Pshycics

OpenStudy (migitmack):

Physics

OpenStudy (anonymous):

\[h(t) = -16t^2 + (Velocity)t + (height)\]

OpenStudy (anonymous):

\[0 = -16t^2 + 25t + 0\] And then you have to factor...hold on a minute

OpenStudy (anonymous):

Ok its easier to understand like this: \[height = -16t^2 + (velocity)t + starting height\] That is a quadratic equation: y = ax^2 + bx + c 0 = -16t^2 + 25t + 0 So you can use the quadratic formula: \[x = (-b \pm \sqrt{b^2 - 4ac} / (2a)\] Let's substitute x for t so that the equation is easier to understand: \[0 = -16x^2 + 25x + 0\] \[y = ax^2 + bx + c\] so a = -16 b = 25 c = 0 Substitute these into the quadratic formula: \[x = -25 \pm \sqrt{25^2 - 4(-16)(0)}\]

OpenStudy (anonymous):

Now solve what's inside the square root sign: 25^2 = 25 * 25 = 625 4*-16*0 = 0 625 - 0 = 625. So now find the sq rt of 625 = 25 \ Now you must solve the formula: x = -25 + 25 /(2(-16)) x = 0 / -32 = 0 x = 0 x = -25 - 25 /(2(-16)) x = -50 / -32 x = 1.5625 *****So it will take approximately 1.5625 seconds for the ball to hit the ground:)

OpenStudy (anonymous):

Hope I helped even though I took so long:D